Fermi National Laboratory


Accelerator Update

Friday June 27
The day shift began with Operations monitoring store 2725, with the Antiproton Source (Pbar) stacking, and with MiniBooNE taking beam.

Switchyard experts began a study period at 8:20 AM.

Operations terminated store 2725 at 8:34 AM.

Two Booster RF stations (BRF13 & 14) tripped off at 8:40 AM due to a vacuum problem. They reset okay.

Operations put the TeV into shot setup at 9:15 AM.

At 11:30 AM, Operations established store 2727 with an initial luminosity of 36.225E30. Pbar resumed stacking.

Linac suffered quadrupole problems that messed up the beam for Booster and MiniBooNE at 2:45 PM. Operations called in Linac and controls experts to help solve the problem.

Operations noticed that Booster seemed to operate better with an RF station (BRF9) turned off. So it was turned off at 3:23 PM. (It was turned back on later.)

A Linac expert conducted fifteen minutes of beam tuning at 5:18 PM.

Saturday June 28
The midnight shift began with Operations monitoring store 2727, with MiniBooNE taking beam, with a Main Injector (MI) expert conducting studies, and with Pbar stacking.

MI studies ended at 1:37 AM.

After Tevatron (TeV) experts finished conducting end-of-store smoothing studies, Operations terminated store 2727 at 6:15 AM.

At 7:09 AM, operators made the TeV ready for access into F sector, to change out some equipment, and so the D-zero (D0) experimenters could access their collision hall.

D0 came out of access at 8:08 AM.

Operation put the TeV into shot setup at 9:45 AM. At 12:16 PM, Operations established store 2728 with an initial luminosity of 35.77E30. Pbar resumed stacking.

The TeV store aborted at 7:20 PM due to a RF power supply problem. Coincidentally, a loud thunderclap occurred as the abort fired.

At 7:31 PM, operators made the TeV ready for a Collider Detector Facility (CDF) and D0 access. D0 began their access at 8:18 PM and CDF began theirs at 8:47 PM.

Booster suffered a vacuum burst at 10:05 PM.

The access ended and a TeV expert began a chromaticity study period at 10:47 PM

Sunday June 29
The midnight shift began with Pbar stacking, with a TeV expert conducting studies, and with MiniBooNE taking beam.

The TeV studies ended at 1:40 AM.

Operations put the TeV into shot setup at 2 AM. At 4 AM, Operations established store 2734 with an initial luminosity of 30.75E30. Pbar resumed stacking.

A Booster power supply (ORBMP) tripped of at 1:17 PM and then repeatedly tripped off due to a LCW flow problem. A Booster expert came in and adjusted a flow switch at 3 PM. This switch will have to be replaced at some later date.

Cryo technicians at CHL noticed some problems with the wet engine for sector D4 at 7:11 PM.

A Booster kicker (MKS08) stopped putting out voltage at 9:05 PM. Operations tuned up the backup kicker (MKS04), which worked for stacking and MiniBooNE, but not for other events. Operations called in a Booster expert.

Operations terminated store 2734 at 10:20 PM.

A MI RF station (MIRF12) tripped off at 11:26 PM.

Monday June 30
The midnight shift began with a Booster expert investigating MKS08 problems, which halted stacking and MiniBooNE beam.

The Booster expert finished his work at 12:23 AM. Operations reestablished MiniBooNE beam at 1:02 AM.

Operations established store 2736 at 2:19 AM. Its initial luminosity was 32.84E30. Pbar resumed stacking.

At 4:37 AM, a Linac RF station (LRF5) tripped. Then at 4:46 AM, a Booster RF station (BRF9) tripped off repeatedly followed by another LRF5 trip. Everything reset okay.

Today's Plan
The plan for today is to stack and store. There will be an eight-hour access starting at 8 AM on Wednesday morning for Recycler BPM work in the Main Injector.

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 6/30/2003   email Fermilab